A box contains 7 green marbles and 11 white marbles. If the first marble chosen was a green marble, what is the probability of choosing, without replacement, another green marble? Express your answer as a fraction or a decimal number rounded to four decimal places.
Answer #1

classmate Date Page 0 seli lef A denote the ои first of marble evend of getting event of getting to find the probasility of e
Answer #2
7-1=6 green marbles that’s left Total 7+11=18 18-1= 17 that’s total left marbles 6/17=0.3529 is the correct answer
answered by: Yr